# Byzantine-Eavesdropper Alliance: How to Achieve Symmetric Privacy in Quantum $X$-Secure $B$-Byzantine $E$-Eavesdropped $U$-Unresponsive $T$-Colluding PIR?

### v1

@@ 8 Proofs
Proof:  First, note that since each symbol and each message are encoded using different random variables, it suffices to prove security for only one symbol. Let the $\ell$th symbol of the $m$th message stored in the $n$th database be denoted by $\tilde{W}_{m,\ell}(n)$, thus
@@
Upon measurement, the user retrieves $N$ symbols of $\bm{h}_{1}$, counting the erasure errors as well. In this paper, we examine the vector $\bm{h}_{1}$ and show that symmetric privacy and eavesdropper security are maintained for whatever $N-U$ or $E\leq N-U$ linear combination that can come out of any measurements.
@@
Proof:  Let =$\mathcal{B}=\{i_{1},\ldots,i_{B}\}$ denote the indices of the Byzantine servers. If the $B$ Byzantine servers are cooperating, they will have the noise that should have been added to the generated answers if they were honest as
@@
$\mathbb{P}(\hat{\bm{Z}}_{\mathcal{B}}=\hat{\bm{z}})$ $=\mathbb{P}(\mathtt{V}(\bm{\alpha}_{\mathcal{B}})_{B\times H+M}\bm{Z}^{\prime}_{[H+M]}+\operatorname{diag}(\bm{\alpha}_{\mathcal{B}})^{H+M}\mathtt{V}(\bm{\alpha}_{\mathcal{B}})_{B\times B}\bm{R^{\prime}}_{[B]}=\hat{\bm{z}})$ (130)
@@
The noise symbols used in protecting the side information are secure against the user for any action performed by the $B$ Byzantine servers.
@@
$=\mathbbm{1}\big{(}H+M>\lfloor\frac{N}{2}\rfloor\big{)}I(\bm{Z}^{\prime}_{\mathcal{L}_{1}}(1);\begin{bmatrix}1&\alpha_{n_{1}}&\ldots&\alpha_{n_{1}}^{H+M-1}\\ \vdots&\vdots&\ddots&\vdots\\ 1&\alpha_{n_{B}}&\ldots&\alpha_{n_{B}}^{H+M-1}\end{bmatrix}\begin{bmatrix}Z^{\prime}_{1}\\ \vdots\\ Z^{\prime}_{H+M}\end{bmatrix})$ (141)
@@
where (145) follows from the data processing inequality and the fact that the eavesdropper can also apply the same algorithm applied by the user to know the location of the Byzantine servers and their transmitted noise symbols, $(a)=0$ since all massages are secure against any $B$ servers, and (149) is due to independence between the masking variables and the messages and due to Lemma 6. $\blacksquare$
@@
Furthermore, $\mathtt{M}^{\prime}_{1}\bm{x}$ hides all the message symbols with $2M+2H+2B-N$ independent random variables from the interference symbols. Thus,
@@
Finally, for the last regime, recall that we can write the answers collectively from the $N-U$ serves as follows
@@
where $\mathtt{M}$ is as defined in (40) or (6.1), is a valid $N$-sum box transfer matrix.

### v2

@@ 8 Proofs
[…]
@@
Upon measurement, the user retrieves $N$ symbols of $\bm{h}_{1}$, counting the erasure errors as well. In this paper, we examine the vector $\bm{h}_{1}$ and show that symmetric privacy and eavesdropper security are maintained for whatever NUN-U or $E\leq N-U$ linear combination that can come out of any measurements.
@@
Proof:  Let $\mathcal{B}=\{i_{1},\ldots,i_{B}\}$ denote the indices of the Byzantine servers. If the $B$ Byzantine servers are cooperating, they will have the noise that should have been added to the generated answers if they were honest as
@@
[…]
@@
The noise symbols used in protecting the interference symbols are secure against the user for any action performed by the $B$ Byzantine servers.
@@
[…]
@@
Remark 16
To make it easier to understand intuitively the proof of Lemma 6: $\mathcal{L}_{1}$ and $\mathcal{L}_{2}$ are the interference terms that are not dropped over the air, i.e., decoded by the user upon measurement. In addition, $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ are the remaining masking terms for the shared randomness between the servers. They are all contaminated by the Byzantine noise, however, since the user decodes the Byzantine noise as explained in Section 7, we drop the noise effect. As a numerical example, if $N=17$, $X=5$, $T=4$, $B=2$, and $E=U=0$, then $\mathcal{L}_{1}=\emptyset$, and $\mathcal{L}_{2}=\{9\}$, i.e., no interference terms are received from the first instance while one interference term from the second instance is decoded by the user, $I_{9}(2)$. On the other hand $\mathcal{H}_{1}=\mathcal{H}_{2}=\{1,2\}$, thus $R_{1}(i)$, $R_{2}(i)$, for $i\in[2]$ are received.
@@
where (145) follows from the data processing inequality and the fact that the eavesdropper can also apply the same algorithm applied by the user to know the location of the Byzantine servers and their transmitted noise symbols, $(a)=0$ since all massages are secure against any $B$ servers, (148) is due to (125) and the independence of the noise variables, and (149) is due to independence between the masking variables and the messages and due to Lemma 6.  $\blacksquare$
Remark 17
To see why it is enough to consider $(\mathtt{M}^{\prime}_{1}\bm{x})_{[N+1,2N-2U]}$ for the answers: Considering that the servers share the quantum state and encoding operation done by them, the answers the user gets will be elements of an orthonormal basis for the whole quantum system of all servers. This orthonormality then translates to perfect distinguishability in the measurements, which means that the measurement results completely describe the quantum answers the user has received. Thus, $(\mathtt{M}^{\prime}_{1}\bm{x})_{[N+1,2N-2U]}$, which is the classical measurement result, is actually a description of the quantum answer the user has received.
@@
Furthermore, $\mathtt{M}^{\prime}_{1}\bm{x}$ hides all the message symbols with 2M+2H+2BN2M+2H+2B-N independent random variables from the interference symbols. Thus,
@@
Finally, for the last regime, recall that we can write the answers collectively from the NUN-U serves as follows
@@
where $\mathtt{M}$ is as defined in (40) or (6.1), is a valid NN-sum box transfer matrix.